OpenStudy (anonymous):

sqrt (x + b) = sqrt (x- b) +2

OpenStudy (callisto):

Is the question (1)\[\sqrt{x+b} = \sqrt{x-b+2}\]Or (2)\[\sqrt{x+b} = \sqrt{x-b}+2\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

How do you think this will be solved?

OpenStudy (callisto):

Sorry I made a serious mistake there. But... still, what do you need to find?

OpenStudy (anonymous):

the value of x

OpenStudy (anonymous):

Is \(b\) constant?

mathslover (mathslover):

Ok ! so first of all @broncos1fan \[\huge{\textbf{Welcome to openstudy}}\] So here i go with the solutions/hints : \[\huge{\sqrt{x+b}=\sqrt{x-b}+2}\] \[\huge{(\sqrt{x+b})^2=[(\sqrt{x-b}+2)]^2}\] \[\huge{x+b=x-b+4+2\sqrt{x-b}(2)}\]

mathslover (mathslover):

Will this work @Limitless

OpenStudy (callisto):

@mathslover Do you know if the asker need to solve b or x? either way, we can only express an unknown in terms of another unknowns..

OpenStudy (callisto):

*needs

OpenStudy (anonymous):

@mathslover, yup. Your last line should read \(x+b=x-b+4+4\sqrt{x-b}.\) @broncos1fan can solve for \(x\) from here.

OpenStudy (anonymous):

b is constant

mathslover (mathslover):

\[\huge{x+b-x+b=4+2\sqrt{x-b}2}\] \[\huge{2b-4=4\sqrt{x-b}}\] \[\huge{\frac{2b-4}{4}=\sqrt{x-b}}\] square both sides ... and get the answer in terms of b for x

OpenStudy (anonymous):

@mathslover, no need to solve it that far.

mathslover (mathslover):

ok! sorry for that

OpenStudy (anonymous):

I believe all that is desired by this exercise is learning the trick of removing radicals.

OpenStudy (anonymous):

@broncos1fan, do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wonderful. Have a good day now.

OpenStudy (anonymous):

thank you you too